OpenStudy (anonymous):

Which expression is equivalent to the expression nine x plus x? nine x x nine x ten x nine plus two x

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

ten x

OpenStudy (accessdenied):

9x + x = x(9 + 1) = 10x ten x

OpenStudy (anonymous):

yea 9x+x=10 x

OpenStudy (chriss):

10x